The length of a given cylindrical wire is increased by 100%.Due to the consequent decrease in diameter the change in the resistance of the wire will be?
1)200%
2)100%
3)50%
4)300%

Grade:

1 Answers

Saurabh Koranglekar
5 years ago
Ans 4- 300%

R = rho*l/A

Volume is constant

l’ = 2l then A’ = A/2

R’ = 4R

therefore change in Resistance is 3*R*100/R = 300%